A job placement director claims that the average starting salary for nurses is P12,000. A sample of 10 nurses has a mean of P11,750 and a standard deviation of 200. Is there enough evidence to reject the director’s claim at alpha = 0.05

1
Expert's answer
2022-05-26T10:57:47-0400

Null hypothesis H0: μ = 12000 (claim)

Alternative hypotheses H1: μ ≠ 12000

Since α = 0.05, d.f. = 9 and the test is a two-tailed test, using t-table, the critical values are

"t_c" = -2.262 and 2.262

"t = \frac{11750 - 12000}{\frac{200}{\sqrt{10}}} = -0.395"

We cannot reject the null hypotheses, because -0.395 > -2.262

There is not enough evidence to reject the claim that the starting salary of nurses is $24000.